Can I get some help with this problem? I’m trying to understand this concept so please help with all work step by step shown. Thank you so much!
2. A drug company is comparing the results of three medications on the severity of bacillary dysentery in 15 Peace Corps volunteers returning to the US (N -15). Below are the averages for these three groups Placebo: xbar 11.74, n 5 Ciprofloxacin: xbara 3.64, n Ofloxacin: xbar 3.48, n-5 5 a. State the null and alternative hypotheses. b. Conduct an ANOVA, showing all your work for each step, by completing the ANOVA table below. You need to show both the work for the answers you will present in the table and then also present the completed table itself. Source Sum of Squares d Mean square F-statistic fect 50.075 Total 65.075 Use a Type I error rate of a (alpha) = 0.05, look up your critical F-value. What do you conclude regarding the null hypothesis? Would the p-value be bigger or smaller than . . lpha? What do you conclude about your research question (use your own words, in everyday language)? Conduct and interpret follow-up t-tests for this ANOVA analysis. 497/assignments/951649
Answer #1

a.
Null hypothesis H0: Mean severity of bacillary dysentery of three medications on volunteers are equal.
Alternative hypothesis Ha: At least one medication has different severity of bacillary dysentery on volunteers.

b.

Source Sum of squares df Mean Square F-statistic
Effect 50.075 2 25.0375 20.03
Error 15 12 1.25
Total 65.075 14

df for Total = N-1 = 15 - 1 = 14

df for Effect = number of groups - 1 = 3 - 1 = 2

df for Error = N - number of groups = 15 - 3 = 12

Sum of squares of error = Sum of squares of Total - Sum of squares of Effect

= 65.075 - 50.075 = 15

Mean Square for each Source = Sum of squares / df for Source

Mean Square for Effect = Sum of squares of Effect / df for Effect = 50.075/2 = 25.0375

Mean Square for Error = Sum of squares of Error / df for Error = 15 / 12 = 1.25

F-statistic = Mean Square for Effect / Mean Square for Error = 25.0375 / 1.25 = 20.03

c.

Numerator df = df for Effect = 2

Denominator df = df for Error = 12

Critical value of F at 0.05 significance level and df = 2, 12 is  3.885

d.

As, the observed F-statistic is greater than the Critical value of F, we reject null hypothesis H0. As, we have rejected H0, p-value should be less than significance level of 0.05.

e.

As, we reject null hypothesis H0, we conclude that there is significant evidence that at least one medication has different severity of bacillary dysentery on volunteers.